OpenStudy (anonymous):

A sequence has its first term equal to 4 and each term of the sequence is obtained by adding 2 to the previous term. If f(n) represents the nth term of the sequence, which of the following recursive functions best defines this sequence? (1 point) f(1) = 2 and f(n) = f(n - 1) + 4; n > 1 f(1) = 4 and f(n) = f(n - 1) + 2n; n > 1 f(1) = 2 and f(n) = f(n - 1) + 4n; n > 1 f(1) = 4 and f(n) = f(n - 1) + 2; n > 1

OpenStudy (anonymous):

this is an A.P first term (a) = 4 common difference (d) = 2 Fn=a+(n-1)d therefore, Fn=4 + (n-1) 2 Fn = 4 +(2n-2) F n = 4+2n-2= 2+2n Fn= 2 (1+n)